# Offline Clustering of Linear Bandits: Unlocking the Power of Clusters in Data-Limited Environments

### v1

@@ Abstract
Contextual linear multi-armed bandits are a learning framework for making a sequence of decisions, e.g., advertising recommendations for a sequence of arriving users. Recent works have shown that clustering these users based on the similarity of their learned preferences can significantly accelerate the learning. However, prior work has primarily focused on the online setting, which requires continually collecting user data, ignoring the offline data widely available in many applications. To tackle these limitations, we study the offline clustering of bandits (Off-ClusBand) problem, which studies how to use the offline dataset to learn cluster properties and improve decision-making across multiple users. The key challenge in Off-ClusBand arises from data insufficiency for users: unlike the online case, in the offline case, we have a fixed, limited dataset to work from and thus must determine whether we have enough data to confidently cluster users together. To address this challenge, we propose two algorithms: Off-C2LUB, which we analytically show performs well for arbitrary amounts of user data, and Off-CLUB, which is prone to bias when data is limited but, given sufficient data, matches a theoretical lower bound that we derive for the offline clustered MAB problem. We experimentally validate these results on both real and synthetic datasets.
@@ 1 Introduction
Contextual bandits (Chu et al., 2011; Agrawal and Devanur, 2016) extend the traditional multi-armed bandit (MAB) framework (Robbins, 1952; Lai and Robbins, 1985; Auer et al., 2002) for optimizing online decision-making to settings where decision rewards depend on given contexts, often through a linear function. Recent works have proposed to accelerate the learning in contextual MAB by taking advantage of the clustered structure in many real-world applications, e.g., social networking (Buccapatnam et al., 2013; Delporte et al., 2013; Bitaghsir et al., 2019) and advertisement recommendation (Ek and Stigsson, 2015; Yan et al., 2022; Aramayo et al., 2023). This framework assumes that a given set of users arrives at the system in sequence; users generally arrive multiple times within the sequence (e.g., users visiting a website multiple times and being shown an ad at each visit). These users can be partitioned into clusters that share the same or similar preferences. The learning agent adaptively clusters users into groups based on their observed responses to the actions taken, leveraging the preference similarities within these groups to improve its decisions. This online and adaptive clustering approach enables more efficient and personalized decision-making, which has been extensively studied across diverse and practical settings (Gentile et al., 2014; Li et al., 2019; Ban et al., 2024). Detailed related works can be found in Appendix 2.
@@
Extending bandit clustering algorithms from the online to the offline setting poses a fundamentally new challenge: unlike in the online setting, where data continually arrives during algorithm execution, the offline setting relies on a fixed offline training dataset. The continual arrival of new data in the online setting allows the algorithm to iteratively refine its understanding of the clustering structure; indeed, over $T$ rounds, only $O(\log(T))$ rounds are typically needed to accurately identify clusters. However, in the offline setting, the number of samples per user is inherently limited. For instance, some users may have sparse historical data due to infrequent activity, or data on certain treatments or actions may be scarce. This limitation creates significant challenges in accurately estimating the preferences of users with limited data and thus in clustering users with similar preferences. We must thus develop criteria to confidently identify similar users, and then account for potential clustering errors—caused by limited or noisy data—in designing and analyzing the bandit algorithm. Unlike standard offline statistical learning (Zhang and Zhang, 2014; Cai and Guo, 2017), where the focus is on estimating parameters with confidence, our setting also requires making structural decisions—such as whether to group users together—based on different amounts of data as well as confidence levels, which fundamentally complicates both the algorithm and its analysis.
@@
To address this challenge, we introduce the Offline Clustering of Bandits (Off-ClusBand) problem. In Off-ClusBand, there are $U$ users represented by the set $\mathcal{U}$, each associated with a preference vector. The users are partitioned into $J$ clusters, with users in the same cluster sharing an identical preference vector. An input dataset $\mathcal{D}$ is provided, containing historical samples of the actions taken and rewards received, assumed to be a linear function of user actions and preference vectors. The objective is to design an algorithm that minimizes the suboptimality gap—the difference between the reward achieved by the algorithm and the optimal reward achievable in hindsight—for any given test user $u_{\text{test}}\in\mathcal{U}$ and an arbitrary action set $A_{\text{test}}$, which may not have been observed in the offline dataset.
Contributions. (i) In Section 3, we formalize the Off-ClusBand problem. We then design the Off-C2LUB algorithm (Algorithm 1) in Section 4, which addresses the challenge of inaccurate clustering due to limited user data by introducing a parameter $\hat{\gamma}$ that thresholds similar users. In Theorem 4.2, we upper bound Off-C2LUB’s suboptimality gap, accounting for two types of error. The homogeneous error decreases as more users sharing the same preference vector of the test user are identified, and this reduction accelerates with an increase in $\hat{\gamma}$ (i.e., the threshold for declaring users to be similar). Conversely, the heterogeneous error, which comes from erroneously treating users with distinct preferences as similar to the test user, may increase with $\hat{\gamma}$. By selecting $\hat{\gamma}\leq\gamma$, where $\gamma$ is the minimum gap between any two users from different clusters, results in zero heterogeneous error but higher homogeneous error; while selecting a large $\hat{\gamma}$ can cause a substantial increase in heterogeneous error but evidently decrease the homogeneous error. We analyze the influence of $\hat{\gamma}$ and provide policies for selecting it under different scenarios, nearly matching the lower bound of the suboptimality gap that we derive in Section 6.
(iii) Finally, in Section 7, we validate our algorithms through both synthetic and real-world experiments. These experiments demonstrate the strong performance for the suboptimality gap of Off-C2LUB across both scenarios with different amounts of data, especially when coupled with our proposed $\hat{\gamma}$-selection policies. Off-C2LUB reduces the suboptimality gap by 50% compared to previous offline clustering algorithms. Moreover, they confirm the effectiveness of Off-CLUB under sufficient data and highlight its limitations when data is insufficient.
2 Related Works
Table 1: Recommended Algorithms Based on Data Conditions
@@ 2 Related Works
Offline Bandits. Offline multi-armed bandits (MAB), introduced by  Shivaswamy and Joachims (2012), are a specialized variant of offline reinforcement learning (RL) that utilize pre-collected data for decision-making based on confidence balls for regression problems (Javanmard and Montanari, 2014; Zhang and Zhang, 2014; Cai and Guo, 2017). These methods have been applied in domains such as robotics  (Kumar et al., 2020), autonomous driving  (Yurtsever et al., 2020), and education  (Singla et al., 2021). Some offline MAB studies consider scenarios where the offline distribution remains consistent with the online rewards  (Bu et al., 2020; Banerjee et al., 2022), while others address settings with reward distribution shifts  (Zhang et al., 2019; Cheung and Lyu, 2024). Research has also explored offline data provided sequentially  (Gur and Momeni, 2022) or in distinct groups  (Bouneffouf et al., 2019; Ye et al., 2020; Tennenholtz et al., 2021). Among these,  Li et al. (2022); Wang et al. (2024) focus on offline contextual linear bandits, which are most closely aligned with our work. Recent works have explored offline RL with homogeneous agents  (Woo et al., 2023, 2024), as well as settings involving heterogeneous agents  (Zhou et al., 2024), where each agent serves a heterogeneous user, or heterogeneous data resources  (Wang et al., ; Shi et al., 2023). Our work addresses offline bandits with multiple data resources for heterogeneous users and introduces a technique to identify and merge similar datasets for more accurate predictions, setting it apart from previous studies.
@@
Online Clustering of Bandits. Online clustering of bandits have been extensively studied across various settings and applications, beginning with Gentile et al. (2014) and including studies linking clusters to items for recommendations (Li et al., 2016a, b; Gentile et al., 2017) and exploring non-linear reward functions (Nguyen et al., 2020; Ban et al., 2024). Distributed bandit clustering has been a focus in  (Korda et al., 2016; Cherkaoui et al., 2023), while extensions to dependent arms were explored in  (Wang et al., 2019). Several works focus on the way to simplify the basic assumptions (Li et al., 2019, 2025). Applications range from recommendation systems  (Yang and Toni, 2018; Yan et al., 2022) to vehicular edge computing and dynamic pricing  (Lin et al., 2022; Miao et al., 2022). Recent studies have also addressed privacy concerns  (Liu et al., 2022), dynamic cluster switching  (Nguyen and Lauw, 2014), unified clustering with non-stationary bandits  (Li et al., 2021), corrupted user detection (Wang et al., 2023) and conversational  (Wu et al., 2021; Li et al., 2023; Dai et al., 2024a, b). In contrast to the online setting, where infinite samples can be drawn from unlimited online data, a key challenge in our setting lies in addressing scenarios with limited offline data. To the best of our knowledge, this is the first work to explore clustering bandits specifically in the offline bandit learning setting.
@@ 3 Problem Formulation
In this section, we introduce the setting for the “Offline Clustering of Bandits (Off-ClusBand)” problem. We consider $U$ users, represented as $\mathcal{U}=\{1,\cdots,U\}$, where each user $u$ is associated with a preference vector $\bm{\theta}_{u}$. To capture the heterogeneity within the user population, users are grouped into $J(\leq U)$ disjoint clusters, with all users in the same cluster sharing an identical preference vector. Precisely, we define $\mathcal{V}=\bigcup_{j=1}^{J}\mathcal{V}(j)$, where $\mathcal{V}(j)$ denotes the set of users within cluster $j$, hence $\mathcal{V}(j)\cap\mathcal{V}(j^{\prime})=\emptyset$ for $j\neq j^{\prime}$. Additionally, $\bm{\theta}_{u}=\bm{\theta}_{u^{\prime}}$ if and only if there exists a $j\in[J]$ such that $\{u,u^{\prime}\}\subseteq\mathcal{V}(j)$. The cluster to which user $u$ belongs is denoted by $j_{u}$, and the shared preference vector for users in cluster $\mathcal{V}(j)$ is denoted as $\bm{\theta}^{j}$. Notably, both the actual partition of clusters and the number of clusters remain unknown to the users. For user $u$, we refer to users in the same cluster as homogeneous users, and those in different clusters as heterogeneous users.
In the offline setting, we are given a dataset $\mathcal{D}=\{\mathcal{D}_{1},\cdots,\mathcal{D}_{U}\}$ comprising offline samples collected from $U$ users. For each user $u$, we observe $N_{u}$ samples, forming $\mathcal{D}_{u}=\{(\bm{a}_{u}^{i},r_{u}^{i})\}_{i=1}^{N_{u}}$, where $\{\bm{a}_{u}^{i}\}_{i=1}^{N_{u}}$ represents the $N_{u}$ given actions, and $\{r_{u}^{i}\}_{i=1}^{N_{u}}$ denotes the corresponding rewards, independently drawn as $r_{u}^{i}\sim R(\bm{a}_{u}^{i})$. Here, $R(\bm{a}_{u}^{i})$ is a 1-subgaussian reward distribution associated with the action $\bm{a}_{u}^{i}$, and $\mathbb{E}[r_{u}^{i}]=\langle\bm{\theta}_{u},\bm{a}_{u}^{i}\rangle$ represents the mean reward for user $u$ under action $\bm{a}_{u}^{i}$. Additionally, we assume that the actions adhere to the following assumption (Gentile et al., 2014, 2017; Blaser et al., 2024).
@@
Each action $\bm{a}$ is independently and identically drawn from a fixed distribution $\rho$ over the set $\{\bm{a}\in\mathbb{R}^{d}:||\bm{a}||_{2}\leq 1\}$. The covariance matrix $\mathbb{E}_{\bm{a}\sim\rho}[\bm{a}\bm{a}^{\top}]$ is full rank with a minimum eigenvalue of $\lambda_{a}>0$.
Let $\pi$ denote an algorithm that selects an action from an action set $A_{\text{test}}\subseteq\mathbb{R}^{d}$ for any given user $u_{\text{test}}\in\mathcal{U}$. The goal in this setting is to design an algorithm $\pi$ that minimizes the suboptimality gap for an incoming user $u_{\text{test}}$ to be served, defined as (Jin et al., 2021; Li et al., 2022):
$\text{SubOpt}(u_{\text{test}},\pi,A_{\text{test}})\coloneqq\langle\bm{\theta}_{u_{\text{test}}},\bm{a}_{\text{test}}^{*}\rangle-\langle\bm{\theta}_{u_{\text{test}}},\bm{a}_{\pi(u_{\text{test}})}\rangle,$
where $\bm{a}_{\text{test}}^{*}$ and $\bm{a}_{\pi(u_{\text{test}})}$ represent the optimal (i.e., reward-maximizing) action in $A_{\text{test}}$ and the action selected by $\pi$ for $u_{\text{test}}$, respectively.
4 Algorithm for Generic: Offline Connection-based Clustering of Bandits
This section introduces our first algorithm: Offline Connection-based Clustering of Bandits Algorithm (Off-C2LUB) and a theoretical upper bound on its suboptimality.

### v2

# Offline Clustering of Linear Bandits: The Power of Clusters under Limited Data

@@ Abstract
Contextual multi-armed bandit is a fundamental learning framework for making a sequence of decisions, e.g., advertising recommendations for a sequence of arriving users. Recent works have shown that clustering these users based on the similarity of their learned preferences can accelerate the learning. However, prior work has primarily focused on the online setting, which requires continually collecting user data, ignoring the offline data widely available in many applications. To tackle these limitations, we study the offline clustering of bandits (Off-ClusBand) problem, which studies how to use the offline dataset to learn cluster properties and improve decision-making. The key challenge in Off-ClusBand arises from data insufficiency for users: unlike the online case where we continually learn from online data, in the offline case, we have a fixed, limited dataset to work from and thus must determine whether we have enough data to confidently cluster users together. To address this challenge, we propose two algorithms: Off-C2LUB, which we show analytically and experimentally outperforms existing methods under limited offline user data, and Off-CLUB, which may incur bias when data is sparse but performs well and nearly matches the lower bound when data is sufficient. We experimentally validate these results on both real and synthetic datasets.
@@ 1 Introduction
Contextual bandits (Chu et al., 2011; Agrawal and Devanur, 2016) extend the fundamental multi-armed bandit (MAB) framework (Robbins, 1952; Lai and Robbins, 1985; Auer et al., 2002) for optimizing online decision-making to settings where rewards depend on given contexts, often through a linear function. Recent works have proposed to accelerate the learning in contextual MAB by taking advantage of the clustered structure in many real-world applications, e.g., social networking (Delporte et al., 2013; Bitaghsir et al., 2019) and advertisement recommendation (Yan et al., 2022; Aramayo et al., 2023). This framework assumes that a given set of users arrives at the system in sequence; users generally arrive multiple times within the sequence (e.g., users visiting a website multiple times and being shown an ad at each visit). These users can be partitioned into clusters that share the same or similar preferences. The learning agent adaptively clusters users into groups based on their observed responses to the actions taken, leveraging the preference similarities within these groups to improve its decisions. This online clustering approach enables more efficient and personalized decision-making, which has been extensively studied across diverse settings (Gentile et al., 2014; Li et al., 2019; Ban et al., 2024). Detailed related works can be found in Section˜2.
@@
Extending bandit clustering algorithms from the online to the offline setting poses a fundamentally new challenge: unlike in the online setting, where data continually arrives during algorithm execution, the offline setting relies on a fixed offline dataset. The continual arrival of new data in the online setting allows the algorithm to iteratively refine its understanding of the clustering structure; indeed, over $T$ rounds, only $O(\log(T))$ rounds are needed to accurately identify clusters. However, in the offline setting, the number of samples per user is inherently limited. For instance, some users may have sparse historical data due to infrequent activity, or data on certain treatments or actions may be scarce. This limitation creates significant challenges in accurately estimating the preferences of users with limited data and thus in clustering users with similar preferences. We thus develop criteria to confidently identify similar users, and then account for potential clustering errors—caused by limited or noisy data—in algorithm design and analysis. Unlike standard offline statistical learning (Zhang and Zhang, 2014; Cai and Guo, 2017; Duan and Wang, 2023), which focuses on parameter estimation, our setting also requires structural decisions, including grouping users and action selections. Moreover, multi-task statistical learning (Duan and Wang, 2023) considers related tasks, but it does not address the challenge of sparse and different amounts of heterogeneous data among users. In contrast, our algorithms leverage these differences to distinguish users and guide clustering.
@@
To address this challenge, we introduce the Offline Clustering of Bandits (Off-ClusBand) problem. In Off-ClusBand, there are $U$ users represented by the set $\mathcal{U}$, each associated with a preference vector. The users are partitioned into $J$ clusters, with all users in a cluster sharing the same preference vector (Gentile et al., 2014; Li and Zhang, 2018; Li et al., 2019; Wang et al., 2025; Li et al., 2025). However, the partition, the number of clusters, and the preference vectors themselves are unknown to the algorithm. An input dataset $\mathcal{D}$ is provided, containing historical samples of the actions taken and rewards received for each user, assumed to be a linear function of user actions and preference vectors. The objective is to identify users with similar preferences to aggregate their offline data, and design an algorithm that minimizes the suboptimality gap—the difference between the reward achieved by the algorithm and the optimal reward achievable in hindsight—for any given test user $u_{\text{test}}\in\mathcal{U}$ and an arbitrary action set $\mathcal{A}_{\text{test}}$, which may not have been observed in the offline dataset.
Table 1: Recommended Algorithms in Different Scenarios ($\gamma$ is the minimum gap between users in different clusters).
@@
Contributions. (i) In Section 3, we formalize the Off-ClusBand problem. We then design the Off-C2LUB algorithm (Algorithm 1) in Section 4, which addresses the challenge of inaccurate clustering due to limited user data by introducing a parameter $\hat{\gamma}$ that thresholds similar users. In Theorem˜4.4, we upper bound Off-C2LUB’s suboptimality gap, accounting for two types of error. The noise decreases as more users’ samples are aggregated, and this reduction accelerates with an increase in $\hat{\gamma}$. Conversely, the bias, which comes from aggregating data from users with distinct preferences, may increase with $\hat{\gamma}$. By selecting $\hat{\gamma}\leq\gamma$, where $\gamma$ is the minimum gap between any two users from different clusters, the algorithm incurs zero bias but higher noise. We analyze the influence of $\hat{\gamma}$ and provide policies for selecting it under different scenarios, nearly matching the lower bound of the suboptimality gap that we derive in Section 5.3.
(iii) Finally, in Section 6, we validate our algorithms through synthetic and real-world experiments. These experiments demonstrate the strong performance of Off-C2LUB with different amounts of data, especially with our proposed $\hat{\gamma}-$selection policies. Off-C2LUB reduces the suboptimality gap by over 60% compared to traditional clustering algorithms. Moreover, they confirm the effectiveness of Off-CLUB under sufficient data and highlight its limitations when data is insufficient.
2 Related Works
@@
Offline Bandits. Offline multi-armed bandits (MAB), introduced by  Shivaswamy and Joachims (2012), are a specialized variant of offline reinforcement learning (RL) that utilize pre-collected data for decision-making based on confidence balls for regression problems (Javanmard and Montanari, 2014; Zhang and Zhang, 2014; Cai and Guo, 2017; Duan and Wang, 2023) where Duan and Wang (2023) also considers parameter estimation with heterogeneous data. These methods have been applied in domains such as robotics  (Kumar et al., 2020), autonomous driving  (Yurtsever et al., 2020), and education  (Singla et al., 2021). Some offline MAB studies consider scenarios where the offline distribution remains consistent with the online rewards  (Bu et al., 2020; Banerjee et al., 2022), while others address settings with reward distribution shifts  (Zhang et al., 2019; Cheung and Lyu, 2024). Research has also explored offline data provided sequentially  (Gur and Momeni, 2022) or in distinct groups  (Bouneffouf et al., 2019; Ye et al., 2020; Tennenholtz et al., 2021). Among these,  Li et al. (2022); Wang et al. (2024) focus on offline contextual linear bandits, which are most closely aligned with our work. Recent works have explored offline RL with homogeneous agents  (Woo et al., 2023, 2024), as well as settings involving heterogeneous agents  (Zhou et al., 2024), where each agent serves a heterogeneous user, or heterogeneous data resources  (Wang et al., ; Shi et al., 2023). Our work addresses offline bandits with multiple data resources for heterogeneous users and introduces a technique to identify and merge similar datasets for more accurate predictions, setting it apart from previous studies.
@@
Online Clustering of Bandits. Online clustering of bandits have been extensively studied across various settings and applications, beginning with Gentile et al. (2014) and including studies linking clusters to items for recommendations (Li et al., 2016a, b; Gentile et al., 2017) and exploring non-linear reward functions (Nguyen et al., 2020; Ban et al., 2024). Distributed bandit clustering has been a focus in  Korda et al. (2016); Cherkaoui et al. (2023), while extensions to dependent arms were explored in  Wang et al. (2019). Several works focus on the way to simplify the basic assumptions (Li et al., 2019, 2025). Applications range from recommendation systems  (Yang and Toni, 2018; Yan et al., 2022) to vehicular edge computing and dynamic pricing  (Lin et al., 2022; Miao et al., 2022). Recent studies have also addressed privacy concerns  (Liu et al., 2022), dynamic cluster switching  (Nguyen and Lauw, 2014), unified clustering with non-stationary bandits  (Li et al., 2021), corrupted user detection (Wang et al., 2023b) and conversational  (Wu et al., 2021; Li et al., 2023; Dai et al., 2024a, b). In contrast to the online setting, where infinite samples can be drawn from unlimited online data, a key challenge in our setting lies in addressing scenarios with limited offline data. To the best of our knowledge, this is the first work to explore clustering bandits specifically in the offline bandit learning setting.
@@ 3 Problem Formulation
In this section, we introduce the setting for the “Offline Clustering of Bandits (Off-ClusBand)” problem. We consider $U$ users, represented as $\mathcal{U}=\{1,\cdots,U\}$, where each user $u$ is associated with a preference vector $\bm{\theta}_{u}$. To capture the heterogeneity within the user population, users are grouped into $J(\leq U)$ disjoint clusters, with all users in the same cluster sharing an identical preference vector. Precisely, we define $\mathcal{U}=\bigcup_{j=1}^{J}\mathcal{V}(j)$, where $\mathcal{V}(j)$ denotes the set of users within cluster $j$, hence $\mathcal{V}(j)\cap\mathcal{V}(j^{\prime})=\emptyset$ for $j\neq j^{\prime}$. Additionally, $\bm{\theta}_{u}=\bm{\theta}_{u^{\prime}}$ if and only if there exists a $j\in[J]$ such that $\{u,u^{\prime}\}\subseteq\mathcal{V}(j)$. The cluster to which user $u$ belongs is denoted by $j_{u}$, and the shared preference vector for users in cluster $\mathcal{V}(j)$ is denoted as $\bm{\theta}^{j}$. Notably, both the actual partition of clusters and the number of clusters remain unknown to the algorithm. For user $u$, we refer to users in the same cluster as homogeneous users, and those in different clusters as heterogeneous users.
In the offline setting, we are given a dataset $\mathcal{D}=\{\mathcal{D}_{1},\cdots,\mathcal{D}_{U}\}$ comprising offline samples collected from $U$ users. For each user $u$, we observe $N_{u}$ samples, forming $\mathcal{D}_{u}=\{(\bm{a}_{u}^{i},r_{u}^{i})\}_{i=1}^{N_{u}}$, where $\{\bm{a}_{u}^{i}\}_{i=1}^{N_{u}}$ represents the $N_{u}$ given actions, and $\{r_{u}^{i}\}_{i=1}^{N_{u}}$ denotes the corresponding rewards, independently drawn as $r_{u}^{i}\sim R(\bm{a}_{u}^{i})$. Here, $R(\bm{a}_{u}^{i})$ is a 1-subgaussian reward distribution associated with the action $\bm{a}_{u}^{i}$, and $\mathbb{E}[r_{u}^{i}]=\langle\bm{\theta}_{u},\bm{a}_{u}^{i}\rangle$ represents the mean reward for user $u$ under action $\bm{a}_{u}^{i}$. We also denote the total number of samples for a set of users $\mathcal{S}$ by $N_{\mathcal{S}}=\sum_{v\in\mathcal{S}}N_{v}$. Additionally, we make the following standard action regularity assumption common in clustering of bandits literature (Gentile et al., 2014; Wang et al., 2023a; Dai et al., 2024a; Wang et al., 2025).
@@
Let $\rho$ be a distribution over $\{\bm{a}\in\mathbb{R}^{d}:\|\bm{a}\|_{2}\leq 1\}$ whose covariance matrix $\mathbb{E}_{\bm{a}\sim\rho}[\bm{a}\bm{a}^{\top}]$ is full rank with minimum eigenvalue $\lambda_{a}>0$. For any fixed unit vector $\bm{\theta}\in\mathbb{R}^{d}$, the random variable $(\bm{\theta}^{\top}\bm{a})^{2}$, with $\bm{a}\sim\rho$, has sub-Gaussian tails with variance upper bounded by $\sigma^{2}$. For each dataset $\mathcal{D}_{u}$, each action $\bm{a}_{u}^{i}$ is selected from a finite candidate set $\mathcal{S}_{u}^{i}$ with size $|\mathcal{S}_{u}^{i}|\leq S$ for any $i\in[N_{u}]$, where the actions in $\mathcal{S}_{u}^{i}$ are independently drawn from $\rho$. Moreover, we assume the smoothed regularity parameter $\tilde{\lambda}_{a}=\int_{0}^{\lambda_{a}}\big(1-e^{-\frac{(\lambda_{a}-x)^{2}}{2\sigma^{2}}}\big)^{S}\mathrm{d}x$ is known to the algorithm.
Remark 3.2 (Discussion on ˜3.1).
˜3.1, adapted from Wang et al. (2023a) and widely used in recent clustering of bandits studies (Dai et al., 2024a; Wang et al., 2025), ensures sufficient coverage of the action space so that information is available along every dimension of the preference vector—a key requirement for distinguishing users across clusters. This assumption is particularly suitable when offline actions are drawn from a finite action space with size bounded by $S$ (e.g. in datasets generated by online logging policies with finite action sets per round). Experiments in Section˜6 and Appendix˜E further validate the effectiveness of our algorithms on offline data generated under different distributions and logging policies. For completeness, we defer discussion of alternative forms of the regularity assumption, such as cases where offline samples are drawn independently from an infinite distribution $\rho$, to Appendix A.
Let $\pi$ denote an algorithm that selects an action from an action set $\mathcal{A}_{\text{test}}\subseteq\mathbb{R}^{d}$ for any given user $u_{\text{test}}\in\mathcal{U}$. The goal in this setting is to design an algorithm $\pi$ that minimizes the suboptimality gap for an incoming user $u_{\text{test}}$ to be served, defined as (Jin et al., 2021; Li et al., 2022):
$\text{SubOpt}(u_{\text{test}},\pi,\mathcal{A}_{\text{test}})\coloneqq\langle\bm{\theta}_{u_{\text{test}}},\bm{a}_{\text{test}}^{*}\rangle-\langle\bm{\theta}_{u_{\text{test}}},\bm{a}_{\pi(u_{\text{test}})}\rangle,$
where $\bm{a}_{\text{test}}^{*}$ and $\bm{a}_{\pi(u_{\text{test}})}$ denote the optimal action in $\mathcal{A}_{\text{test}}$ and the action selected by $\pi$, respectively.
4 Algorithm for General Cases: Off-C2LUB
This section introduces our first algorithm: Offline Connection-based Clustering of Bandits Algorithm (Off-C2LUB) and the theoretical results on its suboptimality.